OpenStudy (anonymous):

Help please =*( Line K is in the rectangular coordinate system. If line K is defined by the equation 3y=2x+6 and line K intersects the x-axis at point (a, b), then what is the value of a?

OpenStudy (anonymous):

answer or step-by-step?

OpenStudy (anonymous):

Okay i know the answer to this i need to know why we multiply the Y by 0 instead of the y-intercept.

Nnesha (nnesha):

y-coordinate = 0 not y-intercept

OpenStudy (anonymous):

ook... don't know why Y gets multiplied by 0 in your given solution... but maybe it was smth like: any point on x axis is (x, 0)

Nnesha (nnesha):

like i said yesterday (a,b) <--- b is y-coordinate not y-intercept (b)

OpenStudy (anonymous):

So why is it 3(0)=2(a)+6 instead of 3=2(a)+0

Nnesha (nnesha):

yep right

OpenStudy (anonymous):

So you need the y-intercept to find the x-intercept?

Nnesha (nnesha):

not the y-intercept y-coordinate that's what you need

OpenStudy (anonymous):

But in this case we multiplied the coordinate by 0 and used the y-intercept to find the solution

Nnesha (nnesha):

 to find x-intercept replace y for 0 always!!